# Forget Vectors at Play: Universal Input Perturbations Driving Machine Unlearning in Image Classification

### v1

@@ 6.2 Experiment Results
Conventionally, unlearning effectiveness is typically measured on the original forget set ($\mathcal{D}_{\mathrm{f}}$). However, with the data perturbation-based forget vector, it is also interesting to investigate its unlearning transferability when applied to a new, previously unseen forget set (denoted as $\mathcal{D_{\mathrm{f}}}^{\prime}$) that share similarities with $\mathcal{D}_{\mathrm{f}}$ and are equally appropriate for unlearning. In the context of class-wise forgetting, we consider $\mathcal{D}_{\mathrm{f}}^{\prime}$ using the testing data from the class targeted for unlearning. In the context of data forgetting, we obtain $\mathcal{D}_{\mathrm{f}}^{\prime}$ by applying the data corruption operation GN1 and ET1 used in Fig. 2 to perturb $\mathcal{D}_{\mathrm{f}}$. Tab. 2 shows the UA of the forget vector applied to the unseen forget set $\mathcal{D}_{\mathrm{f}}^{\prime}$. As observed, the unlearning performance of the forget vector remains effective when evaluated on $\mathcal{D}_{\mathrm{f}}^{\prime}$. Among the model update-based unlearning baselines, current SOTA methods such as SCRUB Kurmanji et al. (2024b) and SalUn Fan et al. (2023) also demonstrate generalization to $\mathcal{D}_{\mathrm{f}}^{\prime}$, compared to simpler MU methods like FT, RL, and GA, which show lower transferability, as evidenced in the class-wise forgetting on (CIFAR-10, ResNet-18).
@@
In Figs. 4 and 5, we explore the impact of the forget vector on unlearning and utility retention through input saliency map, using Grad-CAM (Gradient-weighted Class Activation Mapping) Bengio et al. (2013). Using Grad-CAM, we visualize the salient pixels (i.e., regions most influential to model prediction) for both forget and retain images under different unlearning scenarios: (1) the original model (without unlearning), (2) Retrain, (3) SCRUB-based unlearning, and (4) forget vector-based unlearning. In the first three scenarios, we obtain the input saliency maps on raw forget/retain images without the addition of the forget vector, while in the last scenario, the input saliency is applied to images perturbed by the forget vector. As seen in Fig. 4, only the forget vector and Retrain methods effectively unlearn the target input images, evident from the significant shifts in saliency regions compared to those in the original model. In contrast, the MU baseline SCRUB shows minimal saliency shifts, failing to adequately forget the target data in the last two rows. On the other hand, when evaluating input saliency on retain data in Fig. 5, we observe that the forget vector and other unlearning methods produce input saliency maps similar to those of the original model, indicating retained saliency for non-forgotten images. Moreover, we provide additional visualization results in Appendix D.
@@ Appendix A Generalization of MU to Forget Data Shifts
In Fig. A1, we provide additional evaluations of other approximate unlearning methods, including FT, RL, and GA against Gaussian noise at test time with standard deviations of $0.08$ and $0.2$ Hendrycks & Dietterich (2019), and two types of Elastic transformations with parameters $\left(488,170.8,24.4\right)$ and $\left(488,19.52,48.8\right)$ regarding intensity, smoothing and offset for moderate and high-intensity distortions Hendrycks & Dietterich (2019), as well as a $7$-step PGD attack with perturbation strength $\epsilon=8/255$ Goodfellow et al. (2014). The experiments are conducted on the CIFAR-10 dataset using ResNet-18 for the image classification task. As can be seen, the experimental results presented in Fig. A1 are consistent with the findings in Sec.4 of the main paper, further reinforcing the validity of our conclusions. Specifically, as shown in Fig. A1-(a) and (b), model utility, measured by RA (retain accuracy) and TA (testing accuracy), decreases when external perturbations are applied to the evaluation sets compared to its original performance without perturbations. Meanwhile, Fig. A1-(c) and (d) show that unlearning effectiveness measured by UA (unlearning accuracy) and MIA-Efficacy, remains stable despite the presence of these perturbations on the forget set.
@@ C.2 Run-time Efficiency
FT Warnecke et al. (2021) $4.56$ $2.28$ $2.99$ $2.95$
RL Golatkar et al. (2020) $5.61$ $4.75$ $3.58$ $3.59$
GA Thudi et al. (2022a) $1.16$ $1.15$ $0.30$ $0.31$
NegGrad+ Kurmanji et al. (2024a) $4.72$ $4.22$ $3.21$ $3.17$
SalUn Fan et al. (2023) $2.81$ $3.01$ $2.56$ $2.85$
SCRUB Kurmanji et al. (2024b) $1.23$ $1.53$ $1.62$ $1.55$
@@
To evaluate the computational efficiency in MU, we compare the runtime costs of forget vector calculation with those of other model-based unlearning methods. Following Fan et al. (2023), we adopt the evaluation metric run-time efficiency (RTE), i.e., the computation time of applying an MU method in minutes. Notably, all evaluations are conducted on the same computational environment utilizing an NVIDIA A6000 GPU, using a consistent batch size across all MU methods to maintain fair and reliable comparisons. The corresponding results can be found in Tab. A1. By comparing the training time of model-based methods with our proposed approach, we observe that the forget vector method requires higher computational time than model-based unlearning baselines but remains faster than retraining. This increased time is attributed to the larger number of iterations needed to identify the desired universal input perturbations. However, it is worth noting that the forget vector method optimizes fewer parameters, owing to its input-level design, compared to model-based methods. This trade-off highlights the efficiency of the forget vector in terms of parameter scale despite its higher iteration count.
@@ Appendix D Additional Visualization through An Input Saliency Lens
In Figs. A4 and A5, we provide additional visualization results through input saliency map for different methods against forget images and retain images, respectively, using Grad-CAM (Gradient-weighted Class Activation Mapping) Bengio et al. (2013). Consistent with the main paper, we highlight the salient pixels (i.e., regions most influential to model predictions) under four scenarios: (1) the original model (without unlearning), (2) Retrain, (3) SCRUB-based unlearning, and (4) forget vector-based unlearning. For the first three scenarios, saliency maps are generated on raw forget/retain images, whereas for the last scenario, they are applied to images perturbed by the forget vector.

### v2

@@ 6.2 Experiment Results
Conventionally, unlearning effectiveness is typically measured on the original forget set ($\mathcal{D}_{\mathrm{f}}$). However, with the data perturbation-based forget vector, it is also interesting to investigate its unlearning transferability when applied to a new, previously unseen forget set (denoted as $\mathcal{D_{\mathrm{f}}}^{\prime}$) that share similarities with $\mathcal{D}_{\mathrm{f}}$ and are equally appropriate for unlearning. In the context of class-wise forgetting, we consider $\mathcal{D}_{\mathrm{f}}^{\prime}$ using the testing data from the class targeted for unlearning. In the context of data forgetting, we obtain $\mathcal{D}_{\mathrm{f}}^{\prime}$ by applying the data corruption operation GN1 and ET1 used in Fig. 2 to perturb $\mathcal{D}_{\mathrm{f}}$. Tab. 2 shows the UA of the forget vector applied to the unseen forget set $\mathcal{D}_{\mathrm{f}}^{\prime}$. As observed, the unlearning performance of the forget vector remains effective when evaluated on $\mathcal{D}_{\mathrm{f}}^{\prime}$. Among the model update-based unlearning baselines, current SOTA methods such as SCRUB (Kurmanji et al., 2024b) and SalUn (Fan et al., 2023) also demonstrate generalization to $\mathcal{D}_{\mathrm{f}}^{\prime}$, compared to simpler MU methods like FT, RL, and GA, which show lower transferability, as evidenced in the class-wise forgetting on (CIFAR-10, ResNet-18).
@@
In Figs. 4 and 5, we explore the impact of the forget vector on unlearning and utility retention through input saliency map, using Grad-CAM (Gradient-weighted Class Activation Mapping) (Bengio et al., 2013). Using Grad-CAM, we visualize the salient pixels (i.e., regions most influential to model prediction) for both forget and retain images under different unlearning scenarios: (1) the original model (without unlearning), (2) Retrain, (3) SCRUB-based unlearning, and (4) forget vector-based unlearning. In the first three scenarios, we obtain the input saliency maps on raw forget/retain images without the addition of the forget vector, while in the last scenario, the input saliency is applied to images perturbed by the forget vector. As seen in Fig. 4, only the forget vector and Retrain methods effectively unlearn the target input images, evident from the significant shifts in saliency regions compared to those in the original model. In contrast, the MU baseline SCRUB shows minimal saliency shifts, failing to adequately forget the target data in the last two rows. On the other hand, when evaluating input saliency on retain data in Fig. 5, we observe that the forget vector and other unlearning methods produce input saliency maps similar to those of the original model, indicating retained saliency for non-forgotten images. Moreover, we provide additional visualization results in Appendix D.
@@ Appendix A Generalization of MU to Forget Data Shifts
In Fig. A1, we provide additional evaluations of other approximate unlearning methods, including FT, RL, and GA against Gaussian noise at test time with standard deviations of $0.08$ and $0.2$ (Hendrycks & Dietterich, 2019), and two types of Elastic transformations with parameters $\left(488,170.8,24.4\right)$ and $\left(488,19.52,48.8\right)$ regarding intensity, smoothing and offset for moderate and high-intensity distortions (Hendrycks & Dietterich, 2019), as well as a $7$-step PGD attack with perturbation strength $\epsilon=8/255$ (Goodfellow et al., 2014). The experiments are conducted on the CIFAR-10 dataset using ResNet-18 for the image classification task. As can be seen, the experimental results presented in Fig. A1 are consistent with the findings in Sec.4 of the main paper, further reinforcing the validity of our conclusions. Specifically, as shown in Fig. A1-(a) and (b), model utility, measured by RA (retain accuracy) and TA (testing accuracy), decreases when external perturbations are applied to the evaluation sets compared to its original performance without perturbations. Meanwhile, Fig. A1-(c) and (d) show that unlearning effectiveness measured by UA (unlearning accuracy) and MIA-Efficacy, remains stable despite the presence of these perturbations on the forget set.
@@ C.2 Run-time Efficiency
FT (Warnecke et al., 2021) $4.56$ $2.28$ $2.99$ $2.95$
RL (Golatkar et al., 2020) $5.61$ $4.75$ $3.58$ $3.59$
GA (Thudi et al., 2022a) $1.16$ $1.15$ $0.30$ $0.31$
NegGrad+ (Kurmanji et al., 2024a) $4.72$ $4.22$ $3.21$ $3.17$
SalUn (Fan et al., 2023) $2.81$ $3.01$ $2.56$ $2.85$
SCRUB (Kurmanji et al., 2024b) $1.23$ $1.53$ $1.62$ $1.55$
@@
To evaluate the computational efficiency in MU, we compare the runtime costs of forget vector calculation with those of other model-based unlearning methods. Following (Fan et al., 2023), we adopt the evaluation metric run-time efficiency (RTE), i.e., the computation time of applying an MU method in minutes. Notably, all evaluations are conducted on the same computational environment utilizing an NVIDIA A6000 GPU, using a consistent batch size across all MU methods to maintain fair and reliable comparisons. The corresponding results can be found in Tab. A1. By comparing the training time of model-based methods with our proposed approach, we observe that the forget vector method requires higher computational time than model-based unlearning baselines but remains faster than retraining. This increased time is attributed to the larger number of iterations needed to identify the desired universal input perturbations. However, it is worth noting that the forget vector method optimizes fewer parameters, owing to its input-level design, compared to model-based methods. This trade-off highlights the efficiency of the forget vector in terms of parameter scale despite its higher iteration count.
@@ Appendix D Additional Visualization through An Input Saliency Lens
In Figs. A4 and A5, we provide additional visualization results through input saliency map for different methods against forget images and retain images, respectively, using Grad-CAM (Gradient-weighted Class Activation Mapping) (Bengio et al., 2013). Consistent with the main paper, we highlight the salient pixels (i.e., regions most influential to model predictions) under four scenarios: (1) the original model (without unlearning), (2) Retrain, (3) SCRUB-based unlearning, and (4) forget vector-based unlearning. For the first three scenarios, saliency maps are generated on raw forget/retain images, whereas for the last scenario, they are applied to images perturbed by the forget vector.